HUBBLE SPACE TELESCOPE DAILY REPORT #1527 PERIOD COVERED: 0000Z (UTC) 11/30/95 - 0000Z (UTC) 12/01/95 Daily Status Report as of 335/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.2 Completed WF/PC-2 5957 (Snapshot of 3CR Radio Galaxies) The WF/PC-2 was used to make observations for radio galaxy 3C315.0 = to study the morphology of the radio galaxy. The observations were = executed as scheduled, and no anomalies were noted. 1.3 Completed FOS/BL 5781 (Emission Line Variability in the Lensed = Quasar Q0957+561: Carryover) The Faint Object Spectrograph (Blue Side) was used to make = observations of the lensed quasar Q0957+561 (A and B) to study its = emission line variability. From IUE data it was found that the flux = ratio, A/B varies significantly on a time scale of two weeks. The = acquisition for this proposal failed to fine lock backup on FGS #1 only. = HSTAR 5383 was written. Otherwise, the observations were executed as = scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of WF/PC-2 5973 (The Host Galaxies of Radio = Quite High Redshift Quasi-stellar Objects) The WF/PC-2 was used to make observations of the host galaxy of = radio quite high red shift quasi-stellar object 1207-107. These data will = be compared to existing data sets for radio-loud, high-redshift QSOs and = ground based data on radio-quiet, low-redshift QSOs. The observations = were executed as scheduled, and no anomalies were noted. 1.6 Completed FOS/RD 6167 (FOS Cycle 5: Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to make several = dark calibration images. All exposures were internals. All spectra was = done in the image mode. The observations were completed as planned, and = no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 10 An acquisition at 334/170000Z failed to fine lock backup on FGS #1 = only. HSTAR 5383 was written. Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: As noted in yesterday's report, the battery temperatures continue = to operate at a lower range, with battery #6 reaching -5.3 degrees C. = Omitted was the fact that the lower temperatures are expected since the = primary battery heaters have been disabled. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH